OpenStudy (anonymous):

v^(-0.84)=0.0718, solve for v.

OpenStudy (anonymous):

@whpalmer4 @beccaboo333 @Gabylovesyou

OpenStudy (mathmale):

You might want to approach this by taking the log of both sides. Use rules of logs to solve for v.

OpenStudy (anonymous):

So is it ln(0.0718)/ln(-0.84)?

OpenStudy (mathmale):

v^(-0.84)=0.0718 There's more to it than this: ln(0.0718)/ln(-0.84) If v^(-0.84)=0.0718, take the natural log of both sides. Please show your work so I can undrstand your through processes.

OpenStudy (anonymous):

I honestly don't know how to start.

OpenStudy (mathmale):

\[\ln [v ^{-0.84}]=\ln[0.0718]\] Next, \[\ln [v ^{-0.84}]\]becomes -.084 ln v, so you have\[-0.84 \ln v = \ln 0.0718\]You'll need to solve this first for ln v and next for v itself.